# Task Matters: Knowledge Requirements Shape LLM Responses to Context–Memory Conflict

### v2

@@ Appendix C Task Examples
Input You are a question-answering system that strictly answers questions based only on the given context. Do not use external knowledge or make assumptions beyond what is explicitly stated. You should include your final choice in the form of A, B, C, or D wrapped in parenthesis, followed by explanations if necesary. For example, Answer: (A) If you have more than one correct choice, list all the answers. Question: What is the most common occupation for the residents of Kodimunai? Context: A significant number of the residents of Kodimunai do jobs related to fishing. These jobs includes deep sea fishing, shallow water fishing, fishing from the shore (known as karamadi in the local language), fishing with mechanized boats, exporting fish, etc. Choices: A.Aerospace engineering B.Fishing C.IT, medicine, engineering, trading D.in Answer:
@@
Input "You are a knowledgeable question-answering system. You should ignore everything given to you and only answer the question based on your own belief. You can provide justification if needed. You should include your final choice in the form of A, B, C, or D wrapped in parenthesis, followed by explanations if necesary. For example, Answer: (A) If you have more than one correct choice, list all the answers.Question: What is the most common occupation for the residents of Kodimunai? Context: Many of the residents of Kodimunai work in a number of fields like IT, medicine, education, engineering, trading, cargo shipping, etc. Choices: A.Aerospace engineering B.Fishing C.IT, medicine, engineering, trading D.in Answer:
@@
Input You are a knowledgeable question-answering system. You will be given a context, a question, and a list of choices. Your task is to answer the question using your best possible knowledge. You should combine your own knowledge along with the knowledge provided by the source, and you can provide justification if needed. Note that the provided source is not always reliable. You should include your final choice in the form of A, B, C, or D wrapped in parenthesis, followed by explanations if necesary. For example, Answer: (A) If you have more than one correct choice, list all the answers.Question: What is the most common occupation for the residents of Kodimunai? Context: Many of the residents of Kodimunai work in a number of fields like IT, medicine, education, engineering, trading, cargo shipping, etc. Choices: A.Aerospace engineering B.Fishing C.IT, medicine, engineering, trading D.in Answer:
@@
Input Select the correct answers for the following question based on the given contexts. Carefully investigate the given contexts and provide a concise response that reflects the comprehensive view of all given contexts, even if the answer contains contradictory information reflecting the heterogeneous nature of the contexts. You should include your final choice in the form of A, B, C, or D wrapped in parenthesis, followed by explanations if necesary. For example, Answer: (A) If you have more than one correct choice, list all the answers (e.g. Answer: (BC)). Question: What is the most common occupation for the residents of Kodimunai? Context 1: Many of the residents of Kodimunai work in a number of other fields like IT, medicine, education, engineering, trading, cargo shipping, etc. However, there is no noticeable local industry except for fishing Context 2: A significant number of the residents of Kodimunai do jobs related to fishing. These jobs includes deep sea fishing, shallow water fishing, fishing from the shore (known as karamadi in the local language), fishing with mechanized boats, exporting fish, etc. Choices: A.Aerospace engineering B.Fishing C.IT, medicine, engineering, trading D.in Answer:
@@
An example of each task is included in Figure 12 and Figure 13.
Appendix D Raw Performance
((a)) Contextual Knowledge Task
((b)) Parametric Knowledge Task
((c)) Parametric-Contextual Knowledge Task
Figure 14: F1 score of each model on different task types. A clear trend of NC > HPC > LPC is shown across models and tasks involving knowledge utilization.
Both F1 scores and exact match are measured for CK, PK, PCK tasks. The F1 scores are shown in 14.
Figure 15: Error distribution on RAG task.
@@ Appendix D Raw Performance
Mistral-7B KF 1.6 1.5 1.1 1.0
CK 65.3 46.9 45.3 43.5
PK 62.6 40.5 29.2 34.7
PCK 62.4 31.2 20.8 17.9
RAG 54.4 27.3 18.2 15.7
OLMo2-7B KF 0.0 0.0 0.2 0.2
CK 56.8 52.8 51.0 52.0
PK 55.7 33.8 25.0 26.6
PCK 44.3 22.2 14.8 12.5
RAG 41.5 21.3 14.2 11.5
Qwen2.5-7B KFextract 1.6 1.2 0.9 0.8
CK 78.8 63.0 61.2 56.5
PK 82.8 65.6 53.1 55.5
PCK 83.9 42.2 28.4 24.9
RAG 79.5 40.4 27.5 24.2
Table 1: Performance of models.
@@
The performance of each model on the diagnostic data is shown in Table 1.
D.1 Highly Confident Instances
When querying for the model’s parametric knowledge (parametric knowledge collection in fig.˜2), model responses to queries are collected in a binary stance format (e.g., yes/no). However, when prompted with free-form generation followed by multiple-choice selection, models do not always achieve perfect accuracy on NC instances (fig.˜3). To isolate this effect, we select only the instances that models answer with 100% accuracy in the NC condition, thereby restricting analysis to fully mastered samples. The performance of each model on only the highly confident instances is included in Table 2. The results confirm that while the absolute numbers vary slightly, the overall trends observed in the broader dataset persist.
Appendix E Explanation Generation
When seeing a new context contrary to their knowledge, further explanations are more likely to convince a human, who would iteratively update their mental model with new experiences Vosniadou and Brewer (1992). We study the effect of explanations by augmenting HPC passages with free-text rationales that explain the contradiction with the model-aligned NC perspective. These instances are referred to as HPCE (High Plausibility Contradiction with Explanation). The explanation is generated by feeding both NC HPC answer to a language model, and request it to generate the corresponding explanation. An example of HPCE passage is shown in Figure 16. The prompt used for explanation generation is included below.
Base on the given passage, write a coherent and informative passage that naturally explains why $a^{\texttt{HPC}}$} is the correct explanation or conclusion to the question $q$ instead of $a^{\texttt{NC}}$}. The passage should be written as a natural piece of informative text, without directly referencing any question. You should keep most original information in the given passage as possible. Ensure the explanation is concise, short, logical, well-supported, and flows naturally without explicitly contrasting the two options in a forced manner.
@@ Appendix E Explanation Generation
Appendix F Free Generation Setting
F.1 Evaluator Prompts
@@
We created a free generation setting in §4.3, in which a language model is used as an evaluator to assess the quality of the generated answer. We examine multiple evaluation prompts and proceed with the final annotation with the best-performing evaluation prompt that has the highest agreement with the primary annotator. We follow the design of the evaluator in Hou et al. (2024), made several adjustments to achieve a higher Kohen’s $\kappa$ with human annotators. The final evaluator prompt is included in Figure 18. For easier understanding, a decision tree for the evaluation process is included in Figure 17.
@@
Free Generation Evaluation Prompt 1. Do not fact-check or rely on world knowledge. Assume the reference (correct_answers) is correct. Please evaluate the response to a question under relaxed evaluation, where hallucinations, outdated information are allowed, as long as the primary answer is accurate. Each response is evaluated as correct, partially correct, or incorrect. The list of acceptable answers is separated by "|". You can discuss your reasoning after "comment: ", but you should include your final judgement after "evaluation: ". Decision Rules: * correct * correct means the response accurately matches with all the answers in the given correct answer list, or the correct answer can be inferred from the response. * If more than one correct answer is given, the response can contain contradictory answers that reflect the heterogeneous nature of the context, or the response aggregates the different answers and indicates that each answer is possible. * If only one correct answer is given, the response will be considered as correct as long as the response aligns with the correct answer. * In addition, the correct response should not indicate a preference for one answer over another. * The correct response should not combine two different correct answers without indicating the contradictory nature of these answers. * partially correct * partially correct means the response only matches with one of the answers in the correct answer list; * or the response matches with all the answers in the correct answer list but it indicates a preference for one answer over another; * or the response matches with all the answers in the correct answer list but it contains additional answers that doesnt match with any of the correct answers. * incorrect * incorrect means the response does not match with any of the correct answers in the correct answer list, * or the response merely combines two contradictory answers from the correct answer list and indicates that both of them are possible. Note that for each question, there may be multiple correct answers based on different sources even though these correct answers contradict each other. * For answers that involve names of entities (e.g., people), complete names or commonly recognized names are expected. * Regarding numerical answers, approximate numbers are generally not accepted unless explicitly included in the ground-truth answers. * If the instance has only one correct answer, and the response gives additional answers that are not on the correct answer list, it will be considered as incorrect. The primary or final answers when standing alone must be accurate. Any additional information that is provided must not contradict the primary answers or reshape ones perception of them. We accept responses that contain hallucinated or outdated information that does not significantly impact the primary answers. Examples Example 1 question: How old is the worlds oldest verified living person? correct_answers: 115 years old | 117 years old response: As of today the most up-to-date and relevant information regarding this query is as follows. The oldest verified living person is Maria Branyas Morera, who was born on March 4, 1907, making her 117 years old. comment: The primary answer in the response (117 years old) is partially correct as it only matched with with one of the correct answers (117 years old) and it doesnt mention another answer (115 years old). Thus, the response is partially correct. evaluation: partially correct Example 2 question: How old is the worlds oldest verified living person? correct_answers: 115 years old | 117 years old response: According to the context, the oldest verified living person is Maria Branyas Morera, who is both 115 years old and 117 years old. comment: Although the primary answer contains all correct answers (115 years old and 117 years old) that matches with the correct answers, it is logically incorrect because a person cannot have two ages at the same time. Thus, the response is incorrect. evaluation: incorrect
@@
Free Generation Evaluation Prompt 1. Example 3 question: How old is the worlds oldest verified living person?, correct_answers: 115 years old | 117 years old response: According to the context, one source claims that the oldest verified living person is Maria Branyas Morera, who is 117 years old. However, another source claims that she is 115 years old. comment: The answer contains all correct answers: 115 years old and 117 years old, and it points out that these two answers are from different sources, which is logically possible. Thus, the response is correct. evaluation: correct Example 4 question: How old is the worlds oldest verified living person?, correct_answers: 115 years old | 117 years old response: According to the context, the oldest verified living person is Maria Branyas Morera, who is either 117 years old or 115 years old. comment: The answer contains all correct answers: 115 years old and 117 years old that aggregates the different answers from different sources, which is logically possible. Thus, the response is correct. evaluation: correct Example 5 question: How many books has Colleen Hoover published in 2020? correct_answers: 26 books | 27 response: according to some sources, Colleen Hoover has published 26 books in 2020. However, other sources claim that it has published 27 books in 2020. comment: The response accurately provides two answers (26 books and 27 books) based on different sources. It recognized the two sources that provide different answers. These two answers match with all correct answers (26 books, 27). Thus, the response is correct. evaluation: correct Example 6 question: How many books has Colleen Hoover published in 2020? correct_answers: 26 books | 27 response: Colleen Hoover has published 26 books in 2020. comment: The response only provides one answers (26 books) based on the given context, which matched with one of the answers in the correct answer list. Thus, the response is particially correct. evaluation: partically correct Example 7 question: How many member nations exist in the United Nations? correct_answers: 193 members | 194 member states response: As of my last update in April 2023, there were 191 member nations in the United Nations. Please note that this information may have changed since then, so it is advisable to check the latest information from the United Nations or other reliable sources. comment: The primary answer in the response (191 member nations) does not match with any of the correct answers (193 member states or 194 members). Thus, the response is incorrect. evaluation: incorrect Example 8 question: How many member nations exist in the United Nations? correct_answers: 193 members | 194 member states response: Although some sources claim that there are 193 member nations in the United Nations, the UN is made up of 194 member states. 24 comment: Althought the primary answer mentions two answers (193 members, 104 member states) that match with the correct answers, the response indicates a preference of one answer (194 member states) over the other (193 members). Thus, the response is partically correct. evaluation: partically correct
@@
F.2 Human Annotations
We employ two human annotators from our colleagues without pay to perform the annotation for 50 instances. Both annotators are researchers in natural language processing. Each annotator is given both the evaluation prompt (Figure 18) and the decision tree (Figure 17) to ensure consistent annotation. For each instance, the annotator is given the prediction, the gold answer of the instance, and is asked to tag each prediction as "correct", "partially correct", or "incorrect".
Appendix G License of Artifacts
All license of artifacts used in this work can be found in Table 3.

### v3

@@ Appendix D Task Examples
Input You are a question-answering system that strictly answers questions based only on the given context. Do not use external knowledge or make assumptions beyond what is explicitly stated. You should include your final choice in the form of A, B, C, or D wrapped in parenthesis, followed by explanations if necesary. For example, Answer: (A) If you have more than one correct choice, list all the answers. Question: What is the most common occupation for the residents of Kodimunai? Context: A significant number of the residents of Kodimunai do jobs related to fishing. These jobs includes deep sea fishing, shallow water fishing, fishing from the shore (known as karamadi in the local language), fishing with mechanized boats, exporting fish, etc. Choices: A.Aerospace engineering B.Fishing C.IT, medicine, engineering, trading D.Farming Answer:
@@
Input "You are a knowledgeable question-answering system. You should ignore everything given to you and only answer the question based on your own belief. You can provide justification if needed. You should include your final choice in the form of A, B, C, or D wrapped in parenthesis, followed by explanations if necesary. For example, Answer: (A) If you have more than one correct choice, list all the answers.Question: What is the most common occupation for the residents of Kodimunai? Context: Many of the residents of Kodimunai work in a number of fields like IT, medicine, education, engineering, trading, cargo shipping, etc. Choices: A.Aerospace engineering B.Fishing C.IT, medicine, engineering, trading D.Farming Answer:
@@
Input You are a knowledgeable question-answering system. You will be given a context, a question, and a list of choices. Your task is to answer the question using your best possible knowledge. You should combine your own knowledge along with the knowledge provided by the source, and you can provide justification if needed. Note that the provided source is not always reliable. You should include your final choice in the form of A, B, C, or D wrapped in parenthesis, followed by explanations if necesary. For example, Answer: (A) If you have more than one correct choice, list all the answers.Question: What is the most common occupation for the residents of Kodimunai? Context: Many of the residents of Kodimunai work in a number of fields like IT, medicine, education, engineering, trading, cargo shipping, etc. Choices: A.Aerospace engineering B.Fishing C.IT, medicine, engineering, trading D.Farming Answer:
@@
Input Select the correct answers for the following question based on the given contexts. Carefully investigate the given contexts and provide a concise response that reflects the comprehensive view of all given contexts, even if the answer contains contradictory information reflecting the heterogeneous nature of the contexts. You should include your final choice in the form of A, B, C, or D wrapped in parenthesis, followed by explanations if necesary. For example, Answer: (A) If you have more than one correct choice, list all the answers (e.g. Answer: (BC)). Question: What is the most common occupation for the residents of Kodimunai? Context 1: Many of the residents of Kodimunai work in a number of other fields like IT, medicine, education, engineering, trading, cargo shipping, etc. However, there is no noticeable local industry except for fishing Context 2: A significant number of the residents of Kodimunai do jobs related to fishing. These jobs includes deep sea fishing, shallow water fishing, fishing from the shore (known as karamadi in the local language), fishing with mechanized boats, exporting fish, etc. Choices: A.Aerospace engineering B.Fishing C.IT, medicine, engineering, trading D.Farming Answer:
@@
Examples of each task are provided in Figure˜7 and Figure˜8.
Appendix E Raw Performance
@@
Mistral-7B KFextract 1.6 1.4 0.5 0.4
76.7 77.3 83.4 85.2
CK 65.3 28.5 42.2 38.2
65.3 28.9 42.3 38.5
PK 62.6 18.4 6.4 51.4
62.7 18.6 6.5 51.6
PCK 62.4 0 0 9.4
62.5 31.5 34.2 54.5
RAG 54.4 0.2 0.1 7.9
54.7 41.2 38.1 54.4
OLMo2-7B KFextract 0 0 0.7 0
54.4 66.2 66.4 74.3
CK 56.8 48.9 47.2 55.1
57 48.9 48.5 56
PK 55.7 11.9 7.4 31.2
58.5 14.3 9 32.2
PCK 44.3 0 0 5.7
48.2 39 40.8 50.3
RAG 41.5 1.1 0 3.4
47.8 43.3 42.4 46.3
OLMo2-13B KFextract 1.9 0.2 0.2 0.7
67.4 67.3 74.7 73.4
CK 25.9 16 20.2 4.8
26.4 17 21.3 6
PK 57.5 31.6 18 50.4
57.6 32 18 50.4
PCK 76.1 0 0 6.6
76.1 57.2 50.4 42.7
RAG 82 0.4 0 12.7
82.4 60.9 60.5 65.8
Qwen2.5-7B KFextract 1.6 0.9 0.3 0.5
75.1 71.4 78 80.4
CK 78.8 47.2 57.4 42.6
82.1 50.3 64.4 48.2
PK 82.8 48.5 28.1 62.5
84.4 50.3 31.2 65.8
PCK 83.9 0.5 0.9 14.3
84.4 61.5 60.5 69
RAG 79.5 1.3 1.5 14.3
82.3 60.2 59.8 67.1
Qwen2.5-14B KFextract 1 0.7 0.8 0.4
78.4 79.5 82.8 81.2
CK 87.4 70.2 85.5 41.6
90.5 73.2 90.5 44
PK 76.7 25.7 4.1 48.4
86.9 32 9 55.1
PCK 84.3 2.1 1.8 6.8
89.9 62.7 63.7 61.1
RAG 82.4 10.2 9.9 9.7
87.9 59.1 59.9 61
Table 4: Performance of models. Exact Match rows are unshaded; F1 rows are shaded in numeric columns.
We measure both F1 and exact match of each setting. The full performance of each model on the diagnostic data is shown in Table˜4.
E.1 Highly Confident Instances
@@
When querying for the model’s parametric knowledge (parametric knowledge collection in fig.˜2), model responses to queries are collected in a binary stance format (e.g., yes/no). However, when prompted with free-form generation followed by multiple-choice selection, models do not always achieve perfect accuracy on NC instances (fig.˜3). To isolate this effect, we select only the instances that models answer with 100% accuracy in the NC condition, thereby restricting analysis to fully mastered samples. The performance of each model on only the highly confident instances is included in Table˜5. The results confirm that while the absolute numbers vary slightly, the overall trends observed in the broader dataset persist.
Appendix F Instruction Strength
To further disentangle the role of instruction-following ability from conflict-resolution biases, we conducted an additional experiment varying the strength of instructions with Mistral-7B and OLMo-7B. Specifically, we applied three levels of instruction forcefulness:
  • Strong: “You MUST strictly and exclusively use $\ldots$
  • Neutral: “Answer the question based only on $\ldots$
  • Weak: “Try to answer based on $\ldots$
The average performance on each strength level is shown in Table˜6, and the fine-grained performance on each context and task type is shown in Table˜7 Instruction strength influences absolute performance, but the impact varies across models. Mistral-7B remains relatively stable under different prompt formulations, whereas OLMo-7B shows sharp degradation under strong, restrictive instructions. Despite shifts in absolute scores, the relative patterns reported in §4 remain consistent. Performance differences across tasks and evidence types are preserved, and the ordering of task/evidence effects holds across weak, neutral, and strong prompts.
Model Weak Neutral Strong
Mistral-7B 45.66 46.49 49.48
OLMo2-7B 45.21 44.52 30.83
Table 6: Average performance on each prompt strength.
Task Strength NC HPC HPCE LPC
CK Weak 43.69 57.79 69.34 53.16
Neutral 44.76 57.52 71.52 57.90
Strong 46.39 53.85 70.03 54.20
PK Weak 47.91 18.53 7.95 36.50
Neutral 46.60 15.31 7.60 37.47
Strong 55.06 23.06 8.78 46.98
PCK Weak 44.88 46.45 52.64 57.95
Neutral 41.38 50.58 53.52 58.28
Strong 49.59 48.13 53.59 55.78
RAG Weak 51.23 45.95 45.65 50.93
Neutral 51.50 49.36 49.71 50.80
Strong 57.02 57.21 56.42 55.54
Table 7: Performance of Mistral-7B with different instruction strength. While absolute performance varies, the relative performance relationship still holds.
Appendix G Explanation Generation for HPCE
When encountering context that conflicts with prior knowledge, humans are often more persuaded by additional explanations, which help them iteratively update their mental models with new information Vosniadou and Brewer (1992). To study this effect, we augment HPC passages with free-text rationales that explicitly reconcile the contradiction from the model-aligned NC perspective. We denote these augmented passages as HPCE (High Plausibility Contradiction with Explanation). The explanations are generated by providing both the NC and HPC answers to a language model and prompting it to produce a corresponding rationale. An example of an HPCE passage is shown in Figure˜9, and the full prompt used for explanation generation is provided below.
Based on the given passage, write a coherent and informative passage that naturally explains why $a^{\texttt{HPC}}$} is the correct explanation or conclusion to the question $q$ instead of $a^{\texttt{NC}}$}. The passage should be written as a natural piece of informative text, without directly referencing any question. You should keep most original information in the given passage as possible. Ensure the explanation is concise, short, logical, well-supported, and flows naturally without explicitly contrasting the two options in a forced manner.
@@ Appendix G Explanation Generation for HPCE
Appendix H Free Generation Setting
H.1 Evaluator Prompts
@@
In the free generation setting described in §4.3, we employ a language model as an evaluator to assess the quality of generated answers. We experiment with several evaluation prompts and select the one that yields the highest agreement with human annotations. Following the design of the evaluator proposed in Hou et al. (2024), we introduce modifications to improve reliability, achieving a higher Cohen’s $\kappa$ with human annotators. The final evaluator prompt is provided in Figure˜11, and a decision tree illustrating the evaluation process is shown in Figure˜10.
@@
Free Generation Evaluation Prompt 1. Do not fact-check or rely on world knowledge. Assume the reference (correct_answers) is correct. Please evaluate the response to a question under relaxed evaluation, where hallucinations, outdated information are allowed, as long as the primary answer is accurate. Each response is evaluated as correct, partially correct, or incorrect. The list of acceptable answers is separated by "|". You can discuss your reasoning after "comment: ", but you should include your final judgement after "evaluation: ". Decision Rules: * correct * correct means the response accurately matches with all the answers in the given correct answer list, or the correct answer can be inferred from the response. * If more than one correct answer is given, the response can contain contradictory answers that reflect the heterogeneous nature of the context, or the response aggregates the different answers and indicates that each answer is possible. * If only one correct answer is given, the response will be considered as correct as long as the response aligns with the correct answer. * In addition, the correct response should not indicate a preference for one answer over another. * The correct response should not combine two different correct answers without indicating the contradictory nature of these answers. * partially correct * partially correct means the response only matches with one of the answers in the correct answer list; * or the response matches with all the answers in the correct answer list but it indicates a preference for one answer over another; * or the response matches with all the answers in the correct answer list but it contains additional answers that doesnt match with any of the correct answers. * incorrect * incorrect means the response does not match with any of the correct answers in the correct answer list, * or the response merely combines two contradictory answers from the correct answer list and indicates that both of them are possible. Note that for each question, there may be multiple correct answers based on different sources even though these correct answers contradict each other. * For answers that involve names of entities (e.g., people), complete names or commonly recognized names are expected. * Regarding numerical answers, approximate numbers are generally not accepted unless explicitly included in the ground-truth answers. * If the instance has only one correct answer, and the response gives additional answers that are not on the correct answer list, it will be considered as incorrect. The primary or final answers when standing alone must be accurate. Any additional information that is provided must not contradict the primary answers or reshape ones perception of them. We accept responses that contain hallucinated or outdated information that does not significantly impact the primary answers. Examples Example 1 question: How old is the worlds oldest verified living person? correct_answers: 115 years old | 117 years old response: As of today the most up-to-date and relevant information regarding this query is as follows. The oldest verified living person is Maria Branyas Morera, who was born on March 4, 1907, making her 117 years old. comment: The primary answer in the response (117 years old) is partially correct as it only matched with with one of the correct answers (117 years old) and it doesnt mention another answer (115 years old). Thus, the response is partially correct. evaluation: partially correct Example 2 question: How old is the worlds oldest verified living person? correct_answers: 115 years old | 117 years old response: According to the context, the oldest verified living person is Maria Branyas Morera, who is both 115 years old and 117 years old.
@@
Free Generation Evaluation Prompt 1. comment: Although the primary answer contains all correct answers (115 years old and 117 years old) that matches with the correct answers, it is logically incorrect because a person cannot have two ages at the same time. Thus, the response is incorrect. evaluation: incorrect Example 3 question: How old is the worlds oldest verified living person?, correct_answers: 115 years old | 117 years old response: According to the context, one source claims that the oldest verified living person is Maria Branyas Morera, who is 117 years old. However, another source claims that she is 115 years old. comment: The answer contains all correct answers: 115 years old and 117 years old, and it points out that these two answers are from different sources, which is logically possible. Thus, the response is correct. evaluation: correct Example 4 question: How old is the worlds oldest verified living person?, correct_answers: 115 years old | 117 years old response: According to the context, the oldest verified living person is Maria Branyas Morera, who is either 117 years old or 115 years old. comment: The answer contains all correct answers: 115 years old and 117 years old that aggregates the different answers from different sources, which is logically possible. Thus, the response is correct. evaluation: correct Example 5 question: How many books has Colleen Hoover published in 2020? correct_answers: 26 books | 27 response: according to some sources, Colleen Hoover has published 26 books in 2020. However, other sources claim that it has published 27 books in 2020. comment: The response accurately provides two answers (26 books and 27 books) based on different sources. It recognized the two sources that provide different answers. These two answers match with all correct answers (26 books, 27). Thus, the response is correct. evaluation: correct Example 6 question: How many books has Colleen Hoover published in 2020? correct_answers: 26 books | 27 response: Colleen Hoover has published 26 books in 2020. comment: The response only provides one answers (26 books) based on the given context, which matched with one of the answers in the correct answer list. Thus, the response is particially correct. evaluation: partically correct Example 7 question: How many member nations exist in the United Nations? correct_answers: 193 members | 194 member states response: As of my last update in April 2023, there were 191 member nations in the United Nations. Please note that this information may have changed since then, so it is advisable to check the latest information from the United Nations or other reliable sources. comment: The primary answer in the response (191 member nations) does not match with any of the correct answers (193 member states or 194 members). Thus, the response is incorrect. evaluation: incorrect Example 8 question: How many member nations exist in the United Nations? correct_answers: 193 members | 194 member states response: Although some sources claim that there are 193 member nations in the United Nations, the UN is made up of 194 member states. 24 comment: Althought the primary answer mentions two answers (193 members, 104 member states) that match with the correct answers, the response indicates a preference of one answer (194 member states) over the other (193 members). Thus, the response is partically correct. evaluation: partically correct
@@
H.2 Human Annotations
Two colleagues with expertise in natural language processing served as annotators, compensated at standard research assistant rates. They annotated a sample of 50 instances, each using both the evaluator prompt (Figure˜11) and the decision tree (Figure˜10) to ensure consistency. For each instance, annotators were shown the model prediction alongside the gold answer and asked to label the prediction as correct, partially correct, or incorrect.
Appendix I License of Artifacts
All licenses of artifacts used in this work can be found in Table˜8.